Use the following to answer questions 23 – 25

The company invested $200,000 on October 1, 20XA at 4% simple interest for 2 years (24 months).

23. $______________Determine interest revenue for 20XA    

24. $_______________Determine interest revenue for 20XB                  

25. $_______________Determine interest revenue for 20XC  
Answer #1

23. Interest revenue for 20XA = $200000*4%*3/12 = $2000

24. Interest revenue for 20XB = $200000*4%*1 = $8000

25. Interest revenue for 20XC = $200000*4%*9/12 = $6000